# Causally constrained reduced-order neural models of complex turbulent dynamical systems

### v2

@@ III Causal constraints in practice
In what follows, the six variables $x_{i}$ serve as an effective reduced-order description of an underlying high-dimensional dynamical system, from which we aim to learn a causally constrained neural emulator. The parameters and their physical interpretation are reported in Section IV of the SM. $\xi_{i}(t),i=1,...,6$ is a Gaussian white noise with zero mean, unit variance, and $\sigma$ is the noise amplitude. This stochastic forcing mimics the effect of unresolved fast processes. The model is an ideal first test case for our framework of large-scale, low-frequency climate dynamics for two main reasons. First, Eq. (3) features the abstract form $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{L}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, with energy conserving quadratic nonlinearities such that $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$; a structure that naturally arises when geophysical turbulent flows are projected onto orthogonal bases [63, 64, 65, 66]. Second, this system is strongly mixing, with fast decay of autocorrelations and regular, unimodal probability distributions: such properties are relevant for coarse-grained climate dynamics as found both in models and observational data [32, 6, 100, 68], and underlie the reduced-order modeling of realistic, complex turbulent systems [67]. In this respect, the resulting dynamics differ qualitatively from those of prototypical chaotic models like the Lorenz ’63 system [57], which is weakly mixing, possesses exact symmetries, and exhibits a singular invariant measure [68].
We simulate this model for $T=5\cdot 10^{7}$ time-steps at $dt=0.01$, i.e. $5\cdot 10^{5}$ Model Time Units (MTU). This long integration serves as a stationary run to (i) infer the response operators $R_{k,j}(t)$ from data alone and (ii) to build neural emulators. Two emulators, with and without causal constraints, are then compared in terms of their ability to reproduce both stationary and perturbed statistics. While a comprehensive analysis is provided in the Supplemental Material, the main text highlights the identification of causal links and the ability of the causally constrained emulator to capture time-dependent responses in ensemble mean and variance under strong forcings.
@@
Given a long trajectory $\mathbf{x}(t)=[x_{1},...,x_{6}](t)$ of the CdV model, we infer the direct causal links through the steps proposed in the previous Section. The strategy is highlighted step by step in Figure 1. We infer the response operator at the shortest time scale $R_{k,j}(dt)=-\langle x_{k}(dt)s_{j}(\mathbf{x}(0))\rangle$ via score-matching, consider only the off-diagonal elements, and apply the transformation $\ln|R_{k,j}(dt)|$; see Figure 1(a). We then apply a $k$-means algorithm ($k=2$) to split the large responses from the spurious non-causal ones; see Figure 1(b). This information is used to define the Adjacency matrix in Figure 1(c), thus defining the causal functional form of the system to be imposed in the neural emulator; see 1(d). In Section VII of the SM we also report results for the inference of the response operator $R_{k,j}(t)$ over multiple time lags $t$ using both the score-matching and quasi-Gaussian approximations, further illustrating the advantages of the score-matching approach.
Figure 1: Causal constraining of the CdV model (Eq. (3)) from data through the FDT. Panel (a): The 30 off-diagonal responses $\ln|R_{k,j}(dt)|$ (evaluated at the shortest time scale $t=dt$), flattened and plotted against a dummy index. Panel (b): Unsupervised identification of a response threshold via $k$-means clustering ($k=2$): clusters are colored in black and blue. Panel (c): Causal adjacency matrix: $A_{k,j}=1$ is assigned if its corresponding response value lies in the higher centroid cluster; otherwise $A_{k,j}=0$. Panel (d): “causal” deterministic functional form to be enforced in the emulator loss function in Eq. (10).
@@
We learn two stochastic neural models from the long trajectory $\mathbf{x}(t)=[x_{1},...,x_{6}](t)$ of the CdV model: one with causal constraints and one without. The latter referred to as “vanilla” emulator. The causally constrained emulator leverages the Adjacency matrix $A_{k,j}$ inferred in Figure 1(c) to suppress spurious, non-causal terms through the new loss function in Eq. (10). The set $\mathcal{S}$ of pairs $(k,j)$ to be used in Eq. (10) is given by the zeros of $A_{k,j}$ in Figure 1(c) and equal to $\mathcal{S}=\{(1,2),(1,4),(1,5),(1,6),(2,5),(3,6),(4,1),(5,2),(6,3)\}$. Both neural emulators have been trained with the same hyper-parameters, the only difference being the addition of the regularization term as in Eq. (10) for the causally constrained model.
@@
Given the two emulators defined above, we simulate a long trajectory of $T=10^{7}$ time steps. We then examine two stationary statistics: the invariant probability distribution (PDF) and the autocorrelation functions (ACFs). The analysis is reported in Section VI of the SM: both the vanilla and the causally constrained emulator reproduce these stationary statistics well.
@@
We assess the ability of the emulators to reproduce perturbed statistics in the linear-response regime by analyzing the impulse response operator for the ensemble mean and variance, following [29]. A small impulse perturbation $\delta x_{j}(0)$ is applied at $t=0$, and the response of observables $A(x_{k}(t))=x_{k}(t)$ (mean) and $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$ (variance) is measured, $\bm{\mu}(t)$ representing the time-dependent mean. Here, we perturb $x_{1}(0)$ and report the self-response of $x_{1}(t)$ in mean and variance (Fig. 2). The vanilla emulator reproduces the response operator accurately, as expected given the large training dataset and Markovianity of the system [29]. However, the causally constrained emulator yields a systematically improved representation of both mean and variance responses. This improvement persists for the full response operator $R_{k,j}(t)$ across all variables (see Figures 4–5 of the Supplemental Material). A quantitative summary of this improvement is provided by the total time-dependent MSE, aggregated over all $j$ and $k$, between the emulators and the ground truth, reported in Figure 6 of the SM.
Figure 2: Perturbed statistics: linear regime. Self-response operator of mean and variance for the variable $x_{1}$. Panel (a): Time-dependent, response in ensemble mean of variable $x_{1}(t)$ to an impulse perturbation $\delta x_{1}(0)$ imposed on $x_{1}(0)$ at time $t=0$. Formally: $R_{1,1}(t)=\delta\langle x_{1}(t)\rangle/\delta x_{1}(0)$. Panel (b) same as Panel (a) but for response in ensemble variance. Responses are computed using an ensemble of $N_{e}=10^{6}$ members. The magnitude of the impulse perturbation is set to $10^{-1}\sigma_{1}$, $\sigma_{1}$ being the standard deviation of $x_{1}$. Labels: “Numerical” (numerical model), Model vanilla” (vanilla, unconstrained, emulator), and “Model causal” (causally constrained emulator).
@@
We now probe the nonlinear response regime by applying a large step forcing $\mathbf{F}$ to both the numerical model and the neural emulators. As in the linear response analysis, the forcing is applied to $x_{1}$. The forced dynamics are given by $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, where $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, $\sigma_{1}$ being the std. dev. of the $x_{1}$ variable. The forcing amplitude is intentionally set very high to induce a strongly nonlinear response. We simulate an ensemble of $10^{5}$ forced realizations and measure the time-dependent response of the ensemble mean and variance relative to the unperturbed stationary distribution. The forcing alters the system’s attractor through time-dependent feedbacks across degrees of freedom. Figure 3 shows the resulting self-response of the marginal distribution of $x_{1}(t)$; responses of the full system are reported in Figures 7–8 of the Supplemental Material. The numerical model exhibits a pronounced transient response, relaxing toward a new equilibrium after approximately $40$ model time units. Therefore, the external forcing modifies the characteristics of the attractor in state space. The causally constrained emulator accurately reproduces these nonlinear forced responses, particularly in the ensemble mean. By contrast, the vanilla emulator becomes unstable after a short transient under the same forcing. This behavior is robust: when large step forcings are applied to each degree of freedom, the vanilla emulator becomes unstable in three out of six cases, whereas the causally constrained emulator remains stable for all forcings (see Section VIII of the SM). In general, despite the large forcing amplitude, we observe that the causally constrained emulator provides a good approximation of the forced responses in both ensemble mean and variance.
Figure 3: Perturbed statistics: nonlinear regime. Time-dependent response of the ensemble mean and variance of $x_{1}$ to a strong step function forcing. Forcing: step function $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, applied on the right-hand side of the numerical model and emulators. $\sigma_{1}$ is the standard deviation of $x_{1}$. Responses are computed using an ensemble of $N_{e}=10^{5}$ members. Panel (a): response in ensemble mean. Panel (b): response in ensemble variance. Labels: “Numerical” (numerical model), “Model vanilla” (vanilla, unconstrained, emulator), and “Model causal” (causally constrained emulator).
IV Discussion and conclusions
We introduced a strategy that leverages the Fluctuation–Dissipation Theorem and score matching to (i) identify direct causal links from long trajectories of Markovian physical systems and (ii) impose this causal information as constraints in neural emulators. These constraints improve the ability of reduced-order neural emulators to capture responses to both weak and strong external perturbations, even if trained on unforced data alone. In practice, the choice of a Markovian, coarse-grained representation of an underlying high-dimensional dynamical system is itself a central challenge, as emphasized in classical work [79] and revisited in recent studies [72, 62, 29]. However, given an appropriate choice of coarse-grained variables, the proposed constraints can be readily implemented in general neural network architectures. The framework is modular and can be combined with any data-driven estimator of the FDT operator, including recent proposals such as [61, 25], as well as with alternative causal discovery methods, for example [78]. A natural next step is to extend this framework to observational climate data and to develop causally consistent, reduced-order neural models of the global atmosphere and ocean, including within emerging probabilistic architectures such as stochastic interpolants [18, 107, 77]. This perspective also opens the possibility of combining causal constraints with complementary physical constraints, such as energy conservation [7], to jointly enforce stability and physical fidelity in neural emulators. Additionally, in realistic applications, data are often scarce and partially observed, leading to imperfect causal estimates. An important open direction is to assess how accurate such estimates must be - and whether approximate, quasi-Gaussian FDT operators may already suffice [68] - to construct imperfect emulators that nevertheless exhibit skill in responding to external perturbations. Finally, while motivated by climate dynamics, the proposed strategy applies to neural modeling of general complex turbulent dynamical systems in reduced coordinate spaces. In such settings, causal constraints provide a practical and readily implementable addition to general neural network architectures.
@@ Acknowledgements.
This research was supported by Schmidt Sciences, LLC, through the M2LInES project. Computational resources and support were provided in part by the NYU IT High Performance Computing facilities, services, and staff. F.F. is grateful to Rory Basinski and Matthieu Blanke for many enriching discussions on this subject.
@@
Causally constrained reduced-order neural climate emulators via response theory

### v3

@@ III.1 Stochastic Charney-DeVore model
We integrate Eq. (5) with a RK4 scheme for $T=10^{7}$ time-steps at $dt=0.01$, i.e. $10^{5}$ Model Time Units (MTU). The resulting 6-dimensional trajectory is then treated as a discrete time series $\mathbf{x}_{t}=(x^{(1)}_{t},...,x^{(6)}_{t})$ from which we want to learn the discrete flow map in Eq. (1). The parameters and their physical interpretation are reported in Sec. V of the SM. This stochastic model features fast decay of autocorrelations and regular, unimodal probability distributions: such properties are relevant for coarse-grained climate dynamics as found both in models and observations [26, 4, 89, 60], and underlie the reduced-order modeling of realistic, turbulent systems [59]. The main text highlights the identification of the causal Adjacency matrix and response to weak and strong forcing of variable $x_{1}$; a comprehensive analysis is provided in the SM.
@@
Given a trajectory of the CdV model, we infer the direct causal links. The strategy is summarized in Figure 1. We infer the response operator at the shortest time scale $R^{k,j}_{1}=-\langle x^{(k)}_{1}s_{j}(\mathbf{x}_{0})\rangle$, apply the transformation $\ln|R^{k,j}_{1}|$ and split the large responses from the spurious non-causal ones via k-means clustering (k=2); see Figure 1(a). This information is used to define the Adjacency matrix in Figure 1(b).
Figure 1: Causal constraining of the CdV model (Eq. (5)) from data using the FDT and score matching. Panel (a): The 30 off-diagonal responses $\ln|R^{k,j}_{1}|$, flattened and plotted against a dummy index. Causal links are identified via k-means clustering ($k=2$), shown as filled and open circles. Panel (b): Resulting causal adjacency matrix $A^{k,j}$: $A^{k,j}=1$ (dark) for responses in the cluster with the larger centroid, and $A_{k,j}=0$ (white) otherwise. Example: the first row $A^{1,j}$ implies $x^{(1)}_{t+1}=f_{1}(x^{(1)}_{t},x^{(3)}_{t})$.
@@
We learn two discrete physics-constrained neural models, with and without causal constraints, from the long trajectory $\mathbf{x}_{t}$. We denote these models as “Physics” and “Physics &\& Causal” emulators.
@@
We simulate a long trajectory of $T=10^{7}$ time steps with the neural models. We examine two stationary statistics: the invariant distribution (PDF) and the autocorrelation functions (ACFs). Both emulators faithfully reproduce these stationary statistics; see Sec. VII of the SM.
@@
We analyze the impulse response operator for the ensemble mean and variance. A small impulse perturbation $\delta x^{(j)}_{0}$ is applied at $t=0$, and the response of observables $A(x^{(k)}_{t})=x^{(k)}_{t}$ (mean) and $A(x^{(k)}_{t})=(x^{(k)}_{t}-\mu^{(k)}_{t})^{2}$ (variance) is measured, $\bm{\mu}_{t}$ representing the time-dependent mean (see SM Sec. VI). As a representative example, we perturb $x^{(6)}_{0}$ and report the self-response of $x^{(1)}_{t}$ in mean and variance (Fig. 2(a,c)). Both emulators reproduce the response operator well, with the causally constrained model outperforming the non-causal one. This improvement is robust, as shown in SM Sec. VIII, where we report the time-dependent MSE of the responses $R^{k,j}_{t}$ aggregated over all pairs $j,k$. By contrast, we also trained a vanilla MLP network (no physics and no causal constraints). The vanilla model reproduced well the stationary distribution of the system, but suffered instability even under small perturbations (not shown).
Figure 2: Perturbed statistics. Panel (a) and (c): linear response regime. Panel (a): Time-dependent, response in ensemble mean of variable $x^{(1)}_{t}$ to an impulse perturbation $\delta x^{(6)}_{0}$ imposed on $x^{(6)}_{0}$ at time $t=0$. Formally: $R^{1,6}_{t}=\delta\langle x^{(1)}_{t}\rangle/\delta x^{(6)}_{0}$. Panel (c): same as Panel (a) but for response in ensemble variance. Panel (b) and (d): nonlinear response regime. Panel (b): Time-dependent response of the ensemble mean of $x^{(3)}$ to a step function forcing. Forcing: step function $\mathbf{F}=(0,0,0,0,0,\sigma_{6})$ for $t\geq 0$, applied on the right-hand side of the numerical model and emulators. $\sigma_{6}$ is the standard deviation of $x^{(6)}$. Panel (d): same as Panel (b) but for the ensemble variance. Notably, an unconstrained vanilla MLP reproduces stationary statistics well but becomes unstable even under small perturbations (not shown).
@@
We probe the nonlinear response by applying a large step forcing $\mathbf{F}$ to both the numerical model and the neural emulators. As in the linear analysis, the forcing acts on $x^{(6)}$, $\mathbf{F}=(0,0,0,0,0,\sigma_{6})$ for $t\geq 0$, where $\sigma_{6}$ is the standard deviation of $x^{(6)}$. We simulate an ensemble of $10^{5}$ realizations and measure the response of the ensemble mean and variance relative to the unperturbed stationary distribution. Results are shown in Figure 2(b,d). The numerical model exhibits a pronounced transient before relaxing to a new equilibrium after $\sim 40$ model time units, reflecting a shift of the attractor in state space. Both emulators remain stable under this strong forcing and predict the response well. The physics- and causally-constrained model generally outperformes the physics-only emulator. This improvement is robust as shown in Sec. IX where we report the time-dependent MSE of the responses aggregated over all perturbation experiments.
IV Symmetry-broken Lorenz-96 system
We define a symmetry-broken version of the Lorenz-96 (L96) [51] system as
$\dot{x}_{j}=(x_{j+1}-x_{j-2})x_{j-1}-x_{j}+F_{j}+\sigma\xi_{j}(t),$ (6)
with $j=1,\dots,J$ and periodic boundary conditions $x_{j}=x_{j+J}$. We set $J=20$. The time-indipendent forcing is defined as $F_{j}=F+\eta_{j}$, where $\eta_{j}\sim\mathcal{N}(0,\sigma_{F}^{2})$ with $\sigma_{F}=10$. Because the forcing includes a random component, we impose a minimum value of $4$ to maintain positivity. This site-dependent forcing (ranging from $4$ to $\sim 31$) breaks translational invariance and induces spatially heterogeneous turbulence. Physically, the localized forcing can be interpreted as topographic effects. The stochastic term $\sigma\xi_{j}(t)$, with $\sigma=5$, represents unresolved fast processes. The system in Eq. (6) serves as a prototype of a turbulent dynamical system, with strong mixing, exponential decay of autocorrelations, many unstable directions, and regular distributions even in the deterministic limit [58, 59]. For reproducibility, the specific $F_{j}$ values are reported in SM Sec. X. We simulate a long trajectory of $T=10^{6}$ time steps with $dt=0.01$ ($10^{4}$ model time units). The 20-dimensional trajectory is treated as a discrete time series $\mathbf{x}_{t}=(x^{(1)}_{t},...,x^{(20)}_{t})$ from which we want to learn the discrete flow map in Eq. (1). We infer the causal adjacency matrix following the protocol described in the previous Sections. As shown in Fig. 3, the framework largely recovers the correct causal structure, including the periodic boundary conditions, despite the turbulent nonlinear dynamics. Two false negatives and one false positive are identified (red and blue in Fig. 3). Because the inferred causal structure is imperfect, causal constraints are imposed through the robust loss in Eq. (4).
Figure 3: Same as Figure 1, but for the L96 system defined in Eq. (6). Red squares indicate false negatives: true physical links erroneously labeled as noncausal. The blue square indicates a false positive: a spurious link erroneously labeled as causal. Example: the first row $A^{1,j}$ indicates a functional dependency $\mathbf{x}^{(1)}_{t+1}=f_{1}(x^{(19)}_{t},x^{(20)}_{t},x^{(1)}_{t},x^{(2)}_{t})$
We then train a “Physics” and a “Physics &\& Causal” constrained neural emulators. Both reproduce the invariant density and autocorrelation functions (SM Sec. XA) and accurately capture the linear responses (not shown). Here we probe the nonlinear response regime by applying a Gaussian forcing centered at lattice site $j=10$,
$G_{j}$ $=w\exp\!\left(-\frac{d(j,\mu)^{2}}{2\ell^{2}}\right),$ (7)
$d(j,\mu)$ $=\min(|j-\mu|,\,J-|j-\mu|),$
with $\mu=10$ and $\ell=3$. The distance function $d(j,\mu)$ is needed to account for the periodic boundary conditions. The forcing amplitude is determined by computing the total standard deviation $\sigma_{\text{Tot}}$ (which is on the order of the individual standard deviation $\sigma_{j}$ at each site $x^{(j)}$) and setting $w=10\sigma_{\text{Tot}}$. This constitutes a severe perturbation, an order of magnitude larger than the natural variability of any single degree of freedom. As depicted in Fig. 4(top-left and bottom-left panels), this strong forcing significantly shifts the mean and variance of the attractor. The purely physics-constrained emulator already captures these statistical changes accurately. Augmenting the model with causal constraints yields a slightly improved representation of the response, particularly in the variance. Importantly, we emphasize that given the robust causal penalty in Eq. (4), the causal model maintains this high performance even if the imposed causal priors are imperfect. Finally, we also trained a vanilla MLP without physical or causal constraints: the unconstrained network was more difficult to train, requiring $\sim 10$ times more epochs to achieve a comparable loss. Although this baseline model reproduced the stationary distribution well and remained stable under the large forcing in Eq. (7) (not shown), it exhibited very large errors in the predicted response as also reported in Figure Fig. 4.
Figure 4: Response of the L96 system to a large Gaussian forcing [Eq. (7)] applied and held constant for $t\geq 0$. Top row: ensemble-mean response. Bottom row: ensemble-variance response. Left: ground truth. Second column: difference between the physics-constrained emulator and the truth. Third column: difference between the physics- and causally constrained emulator and the truth. Fourth column: difference between a completely unconstrained (no physics and no causal constraints) model and the truth.
V Discussion and conclusions
We have proposed a framework for physics and causally constrained reduced-order neural modeling of turbulent dynamical systems. The physics-constrained neural models are here formulated as discrete finite-time maps with strictly energy-conserving nonlinearities. This generalizes continuous-time conservation laws into a formulation inherently amenable to coarse-grained, discrete observational data. Building upon this architectural foundation, we demonstrated how causal information, derived via the fluctuation-dissipation theorem (FDT), can be practically enforced to suppress spurious functional dependencies within the neural model. Conservation laws alone do not uniquely determine the system’s dynamics; causal information provides additional constraints on admissible interactions. Crucially, while the energy-conserving architecture is prescribed a priori as a geometric prior, causal constraints must be inferred from data, requiring careful handling of estimation uncertainty. The resulting constrained models yield stable emulators that accurately capture stationary statistics and responses to both weak and strong external forcings, despite being trained exclusively on unforced data. A central challenge for practical data-driven applications remains the construction of a Markovian coarse-grained representation from high-dimensional spatiotemporal data, as emphasized in classical theory [70] and explored in recent work [63, 55, 23]. However, given an approximate Markovian, coarse-grained representation of a spatiotemporal system, informed by physics and domain knowledge, the proposed constrained discrete neural models can be readily adopted. A natural extension of this work is its application to observational climate data to construct reduced-order emulators of the global atmosphere and ocean. More broadly, while our test cases are motivated by climate dynamics, this methodology provides a practical strategy for neural modeling of general complex turbulent systems in reduced coordinate spaces.
@@ Acknowledgements.
This research was supported by Schmidt Sciences, LLC, through the M2LInES project. Computational resources and support were provided in part by the NYU IT High Performance Computing facilities, services, and staff. F.F. is grateful to Rory Basinski, Matthieu Blanke and Andre Souza for enriching discussions on this subject.
@@
Physics and causally constrained discrete-time neural models
of turbulent dynamical systems